OpenStudy (anonymous):

I need some help solving these problems through elimination. 1) 4x+3y=-1 2x+5y=3 2) 12x-5y=9 3x-8y=-18

OpenStudy (amistre64):

Is that the intersection of 4 lines? or 2 seperate questions?

OpenStudy (anonymous):

2 seperate questions. If yiu can guide me through the first, maybe I can get the second?

OpenStudy (amistre64):

elimination just means that we modify one equation thru multiplication in order to find a value that can cancel out a variable in the other equation.

OpenStudy (amistre64):

4x+3y=-1 2x+5y=3 <- we can multiply this by -2 to get: -4x should we?

OpenStudy (anonymous):

Ok..

OpenStudy (amistre64):

4x+3y=-1 (2x+5y=3) (-2) 4x +3y =-1 -4x-10y =-6 <-- now we add them together -------------

OpenStudy (anonymous):

So would y be -1

OpenStudy (amistre64):

4x +3y =-1 -4x-10y =-6 ------------ 0 -7y = -7 <-- now we solve for y y = 1 is correct

OpenStudy (anonymous):

not -1. 1?

OpenStudy (anonymous):

ok. got it

OpenStudy (amistre64):

Since y = 1, we can use that to find "x" in the equations...

OpenStudy (anonymous):

Do I just plug the y into one of the equations? or both?

OpenStudy (amistre64):

either one is fine; they both meet at a specific point if they are not parallel lines; and that point has y=1 in common; so use it in any of them to find the common x value

OpenStudy (anonymous):

Ok, so if I plug it into the first equation I got x=1 is rhat right?

OpenStudy (anonymous):

-1

OpenStudy (amistre64):

-1 is good to me :) this is a generic setup so dont be fooled by the position of the solution in my graph.....

OpenStudy (anonymous):

Oh... thanks, but thankfully my teacher isnt making us graph :)

OpenStudy (amistre64):

You see where the lines cross they have x and y in common right?

OpenStudy (anonymous):

Yes, I do see that.

OpenStudy (amistre64):

the solution to a set of equations is "where the lines have X and Y in common". Or at a single point.

OpenStudy (anonymous):

Thanks for your help. I want to see if I can get the answer to the second one on my own. If I need help will you be around for a few?

OpenStudy (amistre64):

i should be... :)

OpenStudy (anonymous):

Which one should I try to cancel out? Does it matter?

OpenStudy (amistre64):

doesnt matter; pick one and run with it :)

OpenStudy (anonymous):

YAy... I think I got it (3,3) Can you help me with this one or tell me if I'm right. This one is using subsitution. 6x-y=5 y=11x I got x=-5 y=-55 I dont think its right?

OpenStudy (amistre64):

6x - 11x = 5 -5x = 5 x = -1

OpenStudy (amistre64):

since y = "11x" we substitute that value into the other equation: 6x -y = 5 6x - (11x) = 5 ...and solve

OpenStudy (anonymous):

Thank you. I see where I went wrong.